OpenStudy (anonymous):

simplify and put in stranded (2x^2+3x+1) +(2x^2-x+3)

OpenStudy (anonymous):

get rid of the parenthesis and combine like terms... 2x^2 + 2x2 = 3x + -x = 1 + 3 = string those together and you have your simplified form... :)

OpenStudy (anonymous):

string them together with a "+" that is...

OpenStudy (anonymous):

4x^2+2x+4 is your answer by collecting like terms

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

make sure you understand why thats the answer, good uck